OpenStudy (anonymous):

At 10:15 pm, a bust left a bus stop traveling east at 40 mi/h. 2 hours later, another bus left the same stop traveling west at 35 mi/hi. at what time will the buses be 230 miles apart?

OpenStudy (anonymous):

80 + 40*t + 35*t = 230 75t = 230 - 80 t = 150/75 t = 2 hours.

OpenStudy (anonymous):

So, it is four hours after 10:15pm. So, it is at 2:15am.